# Equitable Allocation for Mixtures of Goods and Chores

### v2

@@ 4.3 Type-Normalized Valuations
For any mixed instance with type-normalized valuations and two agents, an EQ1 allocation always exists and can be computed in polynomial time.
@@
The algorithm starts by allocating items from $O^{\pm}$. For every $o\in O^{\pm}$, $o$ is allocated to the agent who values it positively. (Since $o\in O^{\pm}$ and $n=2$, exactly one of the agents values it positively). Once $O^{\pm}$ is allocated entirely, let $v_{1}$ and $v_{2}$ be the utilities derived by the two agents respectively, and say $v_{1}>v_{2}$. Then by type-normalization, there must be enough items in $O^{+}$ such that agent $2$ derives at least $v_{1}-v_{2}$ utility. The algorithm then allocates all such items to agent $2$, starting from its favorite good from the remaining items, until its utility becomes at least as much as agent $1$. Indeed, this partial allocation is EQ1. The remaining instance is the one that contains objective goods and chores. Using Lemma 3.5, we get a complete EQ1 allocation, as desired. ∎
@@
For any mixed instance with type-normalized valuations, two agents, and only subjective items, every PO allocation is EQ. Hence, an EQ allocation always exists and can be computed in polynomial time.
@@ 5 EQ1 and Pareto Optimality
We now discuss equitability in conjunction with PO. Recall that for goods, an EQ1$+$PO allocation may not exist in general but exists for strictly positive valuations; and deciding its existence in general is NP-Hard (Freeman et al., 2019). For chores, such an allocation always exists and admits a pseudo-polynomial time algorithm, but its complexity is an open question (Freeman et al., 2020). We now present our results for mixtures. We first show that an EQ1$+$PO allocation may not exist even for $\{1,-1\}$ type-normalized valuations.
@@
Alice $1$ $1$ $1$ $-1$ $-1$ $-1$
Bob $-1$ $-1$ $-1$ $1$ $1$ $1$
Clara $-1$ $-1$ $-1$ $1$ $1$ $1$
Not only do we encounter the non-existence of EQ1$+$PO allocations, but deciding the existence of such allocations is hard even when the valuations are type-normalized, as shown by the following result.
@@
For any mixed allocation instance, deciding the existence of an EQ1$+$PO allocation is strongly NP-Hard.
@@
We present a reduction from $3$-Partition, known to be strongly NP-hard, where the problem is to decide if a multiset of integers can be partitioned into triplets such that all of them add up to a constant. Formally, the input is a set $S=\{b_{1},b_{2},\ldots b_{3r}\}$; $r\in\mathbb{N}$; and the output is a partition of $S$ into $r$ subsets $\{S_{1},S_{2},\ldots S_{r}\}$ such that $\sum_{b_{i}\in S_{i}}b_{i}=T=\frac{1}{r}\sum_{b_{i}\in S}b_{i}$. Given any instance of $3$-Partition, we construct an instance of allocation problem as follows. We create $r$ set-agents namely $\{a_{1},a_{2},\ldots a_{r}\}$ and one dummy agent, $a_{r+1}$. We create $3r$ set-items namely $\{o_{1},o_{2},\ldots o_{3r}\}$ and two dummy items $\{o_{3r+1},o_{3r+2}\}$. All the set agents value the set items identically at $b_{i}$ and the two dummy items at $-T$ and $-(r-3)T$ respectively. The dummy agent values the set-items $\{o_{1},\ldots o_{3r-1}\}$ at 00, $o_{3r}$ at $-(r-2)T$ and the two dummy items at $T$ and $(r-1)T$ respectively. Note that this is a type-normalized instance such that every agent values all the items together at $2T$. This completes the construction (see Table 4).
@@
Table 4: Reduced instance as in the proof of Theorem 5.2
The equivalence argument is deferred to the appendix. ∎
Towards tractability, we now show that an EQ1$+$PO allocation can be computed efficiently for $\{-1,0,1\}$ valuations (not necessarily normalized), if one exists. We note here that the presence of zeros has been a source of computational challenge for EQ1$+$PO allocations even in the only goods setting (Freeman et al., 2019; Garg and Murhekar, 2024). We briefly describe the algorithm here. We first reduce the partial instance $\mathcal{I}:=O^{\pm}\cup O^{+}$ to a goods-only instance $\mathcal{I^{G}}$ by setting $v_{i}(o)=0$ for all $\{o\in O^{\pm}:v_{i}(o)=-1\}$. We then compute the Nash optimal allocation [^3] Note that the Nash optimal solution maximizes the geometric mean of agents’ utilities. It can be computed for binary valuations in polynomial time (Barman et al., 2018b). $A^{\prime}$ for $\mathcal{I^{G}}$. We extend the partial allocation $A$ to a complete allocation by first allocating the chores to an agent who values them at 00 (if such an agent exists). Otherwise, we use Lemma 3.5 to allocate the remaining chores which are all valued at $-1$ by all the agents. If the completion of $A^{\prime}$ is EQ1$+$PO, Algorithm 2 outputs the complete allocation, otherwise the instance does not admit any EQ1$+$PO allocation.
Theorem 5.3.
For $\{-1,0,1\}$ valuations (even non-normalized), an EQ1$+$PO allocation can be found efficiently, if one exists.
Proof.
We show that Algorithm 2 returns an EQ1$+$PO allocation whenever it exists. Towards correctness, we first establish the following result.
Lemma 5.4.
There is an EQ1$+$PO allocation for the partial instance $\mathcal{I}:=O^{\pm}\cup O^{+}$ if and only if the Nash optimal allocation $A^{\prime}$ for the instance $\mathcal{I^{G}}$ is EQ1.
Proof.
Suppose the Nash optimal allocation $A^{\prime}$ for instance $\mathcal{I^{G}}$ is EQ1. We will show that $A^{\prime}$ is EQ1$+$PO for $\mathcal{I}$ as well. Note that $A^{\prime}$ is EQ1$+$PO for $\mathcal{I^{G}}$ (since Nash satisfies PO). Since $A^{\prime}$ is PO for $\mathcal{I^{G}}$, it never allocates an item to an agent who values it at 00. Therefore, in the instance $\mathcal{I}$, $A^{\prime}$ never allocates any item from $O^{\pm}\cup O^{+}$ to an agent who values it at $-1$ or 00. Hence, all items are allocated to agents who value them the most, and hence $A^{\prime}$ is PO (and EQ1) for $\mathcal{I}$.
On the other hand, suppose there is an EQ1$+$PO allocation $A$ for instance $\mathcal{I}$. If the Nash optimal allocation for $\mathcal{I^{G}}$ is EQ1, we are done. Otherwise, suppose that the Nash optimal allocation for $\mathcal{I^{G}}$ is not EQ1. We will now argue by contradiction that this is not possible. Since $A$ is PO for the instance $\mathcal{I}$, any item $o\in O^{\pm}$ must have been allocated to an agent $a$ who values it at $1$. Indeed, if not, then there is a Pareto improvement by allocating $o$ to $a$, which makes $a$ strictly better off without making any other agent worse off. Likewise, all items $o\in O^{+}$ are allocated to respective agents who value them at $1$. So, we have $\sum_{i\in N}v_{i}(A_{i})=m^{\prime}$ where $m^{\prime}=|O^{\pm}\cup O^{+}|$. Therefore, $A$ is not only PO but also achieves the maximum utilitarian welfare in $\mathcal{I}$. Consider the same allocation $A$ in the instance $\mathcal{I^{G}}$. Since the only modification from $\mathcal{I}$ to $\mathcal{I^{G}}$ is that for the agents who valued items in $O^{\pm}$ at $-1$ in $\mathcal{I}$, now value them at 00 in $\mathcal{I^{G}}$ and everything else remains the same. So, we have $\sum_{i\in N}v_{i}(A_{i})=m^{\prime}$ in $\mathcal{I^{G}}$ as well. Hence, $A$ maximizes the utilitarian welfare in $\mathcal{I^{G}}$ and hence, is PO. Also, $A$ is EQ1 by assumption. Therefore, for the instance $\mathcal{I^{G}}$, $A$ is an EQ1$+$PO allocation. But, Lemma 22 of Freeman et al. (2019) shows that if an instance with binary valuations admits some EQ1 and PO allocation, then every Nash optimal allocation must satisfy EQ1. Therefore, the Nash optimal allocation in $\mathcal{I^{G}}$ satisfies EQ1, which is a contradiction to our assumption. ∎
If $A$ is EQ1$+$PO for $\mathcal{I}$, then we show that the allocation of $O^{-}$ as mentioned above maintains EQ1$+$PO. If $A$ is not EQ1$+$PO, then we show that if the EQ1 violators under $A$ can not be resolved by using the remaining chores, then there is no EQ1$+$PO allocation for the original instance. We defer the details to the appendix. ∎
Algorithm 2 Computing an EQ1$+$PO allocation, if it exists
Input: $n$ agents, $m$ items and $\{-1,0,1\}$ valuations
Output: An EQ1$+$PO allocation, if one exists
1:  $A\leftarrow$ An empty allocation
2:  For $o\in O^{\pm}$ such that $v_{i}(o)=-1$, set $v_{i}(o)=0$.
@@
12:  return  $A$
@@
For any mixed instance with identical valuations, an EQ1$+$PO allocation always exists and admits a polynomial time algorithm.
@@
For $n=2$ and type-normalized $\{-1,0,1\}$ valuations, an EQ1$+$PO always exists and can be computed efficiently.
When we increase the number of agents from $2$ to $3$, an EQ1$+$PO allocation may not exist even with binary type-normalized valuations. Consider $3$ agents and $6$ items such that two of them value the items at $\{1,1,1,0,0,0\}$ and the last agents values at $\{0,0,0,1,1,1\}$. Here, every PO allocation violates EQ1. We note here that the non-existence of EQ1$+$PO allocation for the goods-only setting is due to the presence of items valued at 00 and such allocation always exists for strictly positive valuations (Freeman et al., 2019). But for mixtures, Example 5.1 suggests that even if there are no zero-valued items, an EQ1$+$PO allocation may not exist.
@@ 6 EQX and Social Welfare
In this section, we discuss the computational complexity of finding allocations that maximize the utilitarian welfare (UW) or egalitarian welfare (EW) within the set of EQX allocations for mixtures. Even in the case when all the underlying items are goods, finding a UW or EW allocation within the set of EQ1 allocations (UW/EQ1 or EW/EQ1), for a fixed number of agents, is weakly NP-Hard (Sun et al., 2023). This rules out the possibility of a polynomial-time algorithm for mixed instances. Nonetheless, we present pseudo-polynomial time algorithms for finding such allocations in mixed instances with a constant number of agents. Our algorithmic technique extends those developed by Aziz et al. (2023) to the mixed setting, and relies on dynamic programming.
@@
For any mixed fair division instance with a constant number of agents, computing a Utilitarian or Egalitarian maximal allocation within the set of EQX allocations admits a pseudo-polynomial time algorithm.
@@ 7 Conclusion
Further, we show the hardness of EQ1$+$PO allocations and complement it by efficient algorithms for restricted valuations. Finally, we present a pseudo-polynomial time algorithm for finding welfare-maximizing EQX allocations.
Deciding the existence of EQ1 allocations under normalized/type-normalized valuations for more than two agents and general valuations stands as an interesting open question. An immediate open direction is to resolve the EQ1 existence and computation for instances with valuations $\{-\alpha,0,\beta\}$ where $\alpha\neq\beta$.
Appendix A Omitted Details of Section 3

### v3

@@ 4.3 Type-Normalized Valuations
For any mixed fair division instance with type-normalized valuations and two agents, an EQ1 allocation always exists and can be computed in polynomial time.
@@
The algorithm starts by allocating items from $O^{\pm}$. For every $\in O^{\pm}, oo$ is allocated to the agent who values it positively. (Since $o\in O^{\pm}$ and $n=2$, exactly one of the agents values it positively). Once $O^{\pm}$ is allocated entirely, let $v_{1}$ and $v_{2}$ be the utilities derived by the two agents respectively, and say $v_{1}>v_{2}$. Then by type-normalization, there must be enough items in $O^{+}$ such that agent $2$ derives at least $v_{1}-v_{2}$ utility. The algorithm then allocates all such items to agent $2$, starting from its favorite good from the remaining items, until its utility becomes at least as much as agent $1$. Indeed, this partial allocation is EQ1. The remaining instance is the one that contains objective goods and chores. Using Lemma 3.4, we get a complete EQ1 allocation, as desired. ∎
@@
For any mixed fair division instance with type-normalized valuations, two agents, and only subjective items, every PO allocation is EQ. Hence, an EQ allocation always exists and can be computed in polynomial time.
@@ 5 EQ1 and Pareto Optimality
In this section, we discuss equitability in conjunction with PO. Recall that for goods, an EQ1++PO allocation may not exist in general but exists for strictly positive valuations; and deciding its existence in general is NP-Hard (Freeman et al., 2019a). For chores, such an allocation always exists and admits a pseudo-polynomial time algorithm, but its complexity is an open question (Freeman et al., 2020). We now present our results for mixtures. We first show that an EQ1++PO allocation may not exist even for $\{1,-1\}$ type-normalized valuations, as illustrated by the following example.
@@
Alice 11 11 11 1-1 1-1 1-1
Bob 1-1 1-1 1-1 11 11 11
Clara 1-1 1-1 1-1 11 11 11
Not only do we encounter the non-existence of EQ1++PO allocations, but deciding the existence of such allocations is hard even when the valuations are type-normalized, as shown by the following result.
@@
For any mixed fair division instance, deciding the existence of an EQ1++PO allocation is strongly NP-Hard.
@@
We present a reduction from $3$-Partition, known to be strongly NP-hard, where the problem is to decide if a multiset of integers can be partitioned into triplets such that all of them add up to a constant. Formally, the input is a set $S=\{b_{1},b_{2},\ldots b_{3r}\}$; $r\in\mathbb{N}$; and the output is a partition of $S$ into $r$ subsets $\{S_{1},S_{2},\ldots S_{r}\}$ such that $\sum_{b_{i}\in S_{i}}b_{i}=T=\frac{1}{r}\sum_{b_{i}\in S}b_{i}$. Given any instance of $3$-Partition, we construct an instance of allocation problem as follows. We create $r$ set-agents namely $\{a_{1},a_{2},\ldots a_{r}\}$ and one dummy agent, $a_{r+1}$. We create $3r$ set-items namely $\{o_{1},o_{2},\ldots o_{3r}\}$ and two dummy items $\{o_{3r+1},o_{3r+2}\}$. All the set agents value the set items identically at $b_{i}$ and the two dummy items at T-T and (r3)T-(r-3)T respectively. The dummy agent values the set-items $\{o_{1},\ldots o_{3r-1}\}$ at 0, $o_{3r}$ at (r2)T-(r-2)T and the two dummy items at $T$ and (r1)T(r-1)T respectively. Note that this is a type-normalized instance such that every agent values all the items together at $2T$. This completes the construction (see Table 5).
@@
$a_{r+1}$ $\leavevmode\nobreak\ \ldots\leavevmode\nobreak\ 0$ r+2T-r+2T $T$ (r1)T(r-1)T
Table 5: Reduced instance as in the proof of Theorem 5.2
We now argue the equivalence.
Forward Direction.
Suppose $3$-Partition is a yes-instance and $\{S_{1},S_{2},\ldots S_{r}\}$ is the desired solution. Then, the corresponding EQ1++PO allocation, $A$, can be constructed as follows. For $a_{i}:i\in[r]$, we set $A(a_{i})=\{o_{j}:o_{j}\in S_{i}\}$. Finally, $A(a_{r+1})=\{o_{3r+1},o_{3r+2}\}$. It is easy to see that $A$ is an EQ1 allocation since $v_{i}(A_{i}))=T$ for all $a_{i}:i\in[r]$ and $v_{i}(A_{i})\geq v_{r+1}(A(a_{r+1})\setminus\{o_{3r+2}\})=T$. Also, $A$ is PO since it is a welfare-maximizing allocation where each item is assigned to an agent who values it the most.
Reverse Direction.
Suppose that $A$ is an EQ1++PO allocation. Then, both the dummy items $\{o_{3r+1},o_{3r+2}\}$ must be allocated to the dummy agent $a_{r+1}$, who is the only agent who values both the items positively. (Since any set-agent derives a negative utility from these items, allocating them to it violates PO). Since $A$ is also EQ1, this forces every set-agent to derive the utility of at least $T$ under $A$ so that when it removes the item $\{o_{3r+2}\}$ from $A(a_{r+1})$, EQ1 is preserved. ∎
Towards tractability, we now show that an EQ1++PO allocation can be computed efficiently for $\{-1,0,1\}$ valuations (not necessarily normalized), if one exists. It is relevant to note here that the presence of zeros has been a source of computational challenge for EQ1++PO allocations even in the only goods setting (Freeman et al., 2019a). We briefly describe the algorithm here. We first reduce the partial instance $\mathcal{I}:=O^{\pm}\cup O^{+}$ to a goods-only instance $\mathcal{I^{G}}$ by setting $v_{i}(o)=0$ for all $\{o\in O^{\pm}:v_{i}(o)=-1\}$. We then compute the Nash optimal allocation [^4] Note that the Nash optimal solution maximizes the geometric mean of agents’ utilities. It can be computed for binary valuations in polynomial time (Barman et al., 2018b). $A^{\prime}$ for $\mathcal{I^{G}}$. We extend the partial allocation $A$ to a complete allocation by first allocating the chores to an agent who values them at 0 (if such an agent exists). Otherwise, we use Lemma 3.4 to allocate the remaining chores which are all valued at 1-1 by all the agents. If the completion of $A^{\prime}$ is EQ1++PO, Algorithm 2 outputs the complete allocation, otherwise the instance does not admit any EQ1++PO allocation.
Theorem 5.3.
For any mixed fair division instance with symmetric tri-valued (even non-normalized) valuations ($\{-1,0,1\}$), an EQ1++PO allocation can be found efficiently, if one exists.
Proof.
We show that Algorithm 2 returns an EQ1++PO allocation whenever it exists. Towards correctness, we first establish the following result.
Lemma 5.4.
There is an EQ1++PO allocation for the partial instance $\mathcal{I}:=O^{\pm}\cup O^{+}$ if and only if the Nash optimal allocation $A^{\prime}$ for the instance $\mathcal{I^{G}}$ is EQ1.
Proof.
Suppose the Nash optimal allocation $A^{\prime}$ for instance $\mathcal{I^{G}}$ is EQ1. We will show that $A^{\prime}$ is EQ1++PO for $\mathcal{I}$ as well. Note that $A^{\prime}$ is EQ1++PO for $\mathcal{I^{G}}$ (since Nash satisfies PO). Since $A^{\prime}$ is PO for $\mathcal{I^{G}}$, it never allocates an item to an agent who values it at 0. Therefore, in the instance $\mathcal{I}$, $A^{\prime}$ never allocates any item from $O^{\pm}\cup O^{+}$ to an agent who values it at 1-1 or 0. Hence, all items are allocated to agents who value them the most, and hence $A^{\prime}$ is PO (and EQ1) for $\mathcal{I}$.
On the other hand, suppose there is an EQ1++PO allocation $A$ for instance $\mathcal{I}$. If the Nash optimal allocation for $\mathcal{I^{G}}$ is EQ1, we are done. Otherwise, suppose that the Nash optimal allocation for $\mathcal{I^{G}}$ is not EQ1. We will now argue by contradiction that this is not possible. Since $A$ is PO for the instance $\mathcal{I}$, any item $o\in O^{\pm}$ must have been allocated to an agent $a$ who values it at $1$. Indeed, if not, then there is a Pareto improvement by allocating $o$ to $a$, which makes $a$ strictly better off without making any other agent worse off. Likewise, all items $o\in O^{+}$ are allocated to respective agents who value them at $1$. So, we have $\sum_{i\in N}v_{i}(A_{i})=m^{\prime}$ where $m^{\prime}=|O^{\pm}\cup O^{+}|$. Therefore, $A$ is not only PO but also achieves the maximum utilitarian welfare in $\mathcal{I}$. Consider the same allocation $A$ in the instance $\mathcal{I^{G}}$. Since the only modification from $\mathcal{I}$ to $\mathcal{I^{G}}$ is that for the agents who valued items in $O^{\pm}$ at 1-1 in $\mathcal{I}$, now value them at 0 in $\mathcal{I^{G}}$ and everything else remains the same. So, we have $\sum_{i\in N}v_{i}(A_{i})=m^{\prime}$ in $\mathcal{I^{G}}$ as well. Therefore, $A$ maximizes the utilitarian welfare in $\mathcal{I^{G}}$ and hence, is PO. Also, $A$ is EQ1 by assumption. Therefore, for the instance $\mathcal{I^{G}}$, $A$ is an EQ1++PO allocation. But, the following result by Freeman et al. (2019b) shows that if an instance with binary valuations admits some EQ1 and PO allocation, then every Nash optimal allocation must satisfy EQ1.
Lemma 5.5 (Lemma 22, (Freeman et al., 2019b)).
Let $\mathcal{I}$ be a fair division instance with binary valuations. If $\mathcal{I}$ admits some EQ1 and PO allocation, then every Nash optimal allocation must satisfy EQ1.
Therefore, the Nash optimal allocation in $\mathcal{I^{G}}$ satisfies EQ1, which is a contradiction to our assumption. ∎
If $A$ is EQ1++PO for $\mathcal{I}$, then we show that the allocation of $O^{-}$ as mentioned above maintains EQ1++PO. If $A$ is not EQ1++PO, then we show that if the EQ1 violators under $A$ can not be resolved by using the remaining chores, then there is no EQ1++PO allocation for the original instance. We defer the details to the appendix. ∎
Algorithm 2 Computing an EQ1++PO allocation, if it exists
Input: $n$ agents, $m$ items and $\{-1,0,1\}$ valuations
Output: An EQ1++PO allocation, if one exists
@@
2: For $o\in O^{\pm}$ such that $v_{i}(o)=-1$, set $v_{i}(o)=0$
@@
12:return  AA
We now show that EQ1++PO allocation always exists for any number of agents with identical valuations and can be computed efficiently. We also get existence and efficient computation for two agents and $\{-1,0,1\}$ type-normalized valuations.
@@
For any mixed fair division instance with identical valuations, an EQ1++PO allocation always exists and admits a polynomial time algorithm.
@@
For any mixed fair division instance with two agents and type-normalized $\{-1,0,1\}$ valuations, an EQ1++PO always exists and can be computed efficiently.
When we increase the number of agents from $2$ to $3$, an EQ1++PO allocation may not exist even with binary type-normalized valuations. Consider $3$ agents and $6$ items such that two of them value the items at $\{1,1,1,0,0,0\}$ and the last agents values at $\{0,0,0,1,1,1\}$. Here, every PO allocation violates EQ1. We note here that the non-existence of EQ1++PO allocation for the goods-only setting is due to the presence of items valued at 0 and such allocation always exists for strictly positive valuations (Freeman et al., 2019a). But for mixtures, Example 5.1 suggests that even if there are no zero-valued items, an EQ1++PO allocation may not exist.
@@ 6 EQX and Social Welfare
In this section, we discuss the computational complexity of finding allocations that maximize the utilitarian welfare (UW) or egalitarian welfare (EW) within the set of EQX allocations for mixtures (whenever such allocations exist). Even in the case when all the underlying items are goods, finding a UW or EW allocation within the set of EQ1 allocations (UW/EQ1 or EW/EQ1), for a fixed number of agents, is weakly NP-Hard (Sun et al., 2023). This rules out the possibility of a polynomial-time algorithm for mixed instances. Nonetheless, we present pseudo-polynomial time algorithms for finding such allocations in mixed instances with a constant number of agents. Our algorithmic technique extends those developed by (Aziz et al., 2023a) to the mixed setting, and relies on dynamic programming.
@@
For any mixed fair division instance with a constant number of agents, computing a Utilitarian or Egalitarian maximal allocation within the set of EQX allocations (whenever they exist) admits a pseudo-polynomial time algorithm.
@@ 7 Conclusion
Further, we show the hardness of EQ1++PO allocations and complement it by efficient algorithms for restricted valuations. Finally, we present a pseudo-polynomial time algorithm for finding welfare-maximizing EQX allocations.
Deciding the existence of EQ1 allocations under normalized/type-normalized valuations for more than two agents and general valuations stands as an interesting open question. An immediate open direction is to resolve the EQ1 existence and computation for instances with valuations ternary valuations ($\{-\alpha,0,\beta\}$ where $\alpha\neq\beta$). Such valuations have been considered in the context of EF1 (Bhaskar et al., 2024) and EF1+PO (Aleksandrov and Walsh, 2020).
Appendix A Additional Related Work
As mentioned previously, Aleksandrov [Aleksandrov, 2020] termed equitability as jealousy-freeness (JF), and defined its up-to-one-style-relaxations JFX and JF1 wherein, in addition to requiring that the hypothetical removals are restricted to non-zero valued items, these definitions also allow hypothetical addition of an item to a bundle. Aleksandrov [Aleksandrov, 2023] segregated the notions where addition/duplication is allowed and renamed these notions to DJFX and DJF1 (that is, JFX and JF1 up to removal or duplication). Their JFX definition is analogous to our EQX definition, albeit they restrict the removals to non-zero valued items. However, we note that their Theorem 3 does not hold true. In particular, Algorithm $1$ fails to achieve JFX. Reason being, Algorithm $1$ proceeds in the following greedy manner: until all goods are allocated, it picks the least utility agent and gives them their most preferred remaining good. Then, until all chores are allocated, it picks the greatest utility agent and gives them their least preferred remaining chore. Our Example 3.3 shows that such an algorithm does not produce an EQX allocation (which coincides with JFX allocation in this case) even with identical valuations and two agents. Removal of any non-zero valued chore from the bundle of Alice (who is jealous of Bob) works to get rid of the inequitability, however, removal of any non-zero valued good from Bob’s bundle still results in Bob being the richer agent.
Appendix B Omitted Details of Section 3